If p = dy/dx, show that d^2y/ dx^2 = p(dp/dy).
Hence find the solution y = f(x) of the differential equation y(d^2y/dx^2) = 2(dy/dx) + (dy/dx)^2.
1
Expert's answer
2021-01-26T03:20:37-0500
"\dfrac{d^2y}{dx^2}=\dfrac{d}{dx}(\dfrac{dy}{dx})=\dfrac{dp}{dx}=\dfrac{dp}{dy}(\dfrac{dy}{dx})=p\dfrac{dp}{dy}"

"p=\dfrac{dy}{dx}"

Then


"y(p\dfrac{dp}{dy})=2p+p^2"

"p=0\ or\ y\dfrac{dp}{dy}=p+2"

"p=0=>\dfrac{dy}{dx}=0=>y=C_1"


"\dfrac{dp}{p+2}=\dfrac{dy}{y}"

"\int\dfrac{dp}{p+2}=\int\dfrac{dy}{y}"

"p+2=C_2y"

"\dfrac{dy}{dx}=C_2y-2"

"\dfrac{dy}{C_2y-2}=dx"

"\dfrac{1}{C_2}\ln|C_2y-2|=x+\dfrac{1}{C_2}\ln C_3"

"C_2y-2=C_3e^{C_2x}"

"y=\dfrac{2}{C_2}+C_4e^{C_2x}, C_2\not=0"

"y=-2x+C_5, C_2=0"

"y=C_1"

"y=-2x+C_5"

"y=\dfrac{2}{C_2}+C_4e^{C_2x}, C_2\not=0"